# Fractional Chern insulator candidate in twisted bilayer checkboard lattice

### v1

@@ VI Conclusion and discussion
In this manuscript, we study a putative $C=2$ FCI phase in $\nu=1/5,\kappa=0$ MATBCB. In the single particle level, we obtain the lowest topological flat bands with $C=\pm 2$. The band width is about $1.5\times 10^{-3}{\rm meV}$ at the first magic angle $\phi=1.608^{\circ}$. The almost uniform Berry curvature and tiny trace condition violation indicate the ideal quantum geometry condition to host FCI phase. In the many-body level, we first present the many body energy spectrum and spectrum flow for projected Coulomb interaction model of MATBCB. We find there is topological ground state manifold candidate with GSD=10. Next we examine the PES quasi-hole counting of the ground state manifold candidate. We find the counting meets the $(1,5)_{C=2}$ GPP as Halperin spin singlet. This counting is a smoking gun evidence of $C=2$ FCI. When the parameters are not ideal, such as the chiral ratio deviates from zero, there will be a phase transition which breaks the FCI phase. We estimate that $\kappa_{c}\approx 0.7050$. More comprehensive research on the FCI relevant phase transition will be left for future work.
However, one should still keep in mind that the results and arguments above are all built on the assumption that the translation symmetry is preserved. When the translation symmetry is broken, the internal SU(C) color symmetry can be entangled with spatial symmetry. Especially, the SU(C) symmetry and spatial translation symmetry will combine to form a generalized (pseudo) magnetic translation symmetry [33]. In that case, we conjecture that the SU(C) color symmetry will also be broken accompanied with the translation symmetry broken. An example is the $\nu=1/6,C=2$ flat band system [65]. At that situation, Halperin color singlet quasi-hole counting ansatz based $(k,r)_{C}$ GPP may be violated. A more general pseudo potential method under $n_{x}-k_{y}$ basis has to be used [33, 8, 94]. Moreover, under $n_{x}-k_{y}$ basis it is convenient to simulate fractional filling system with numerical methods like DMRG [63, 62], which can handle larger size compared with ED. Compared with analytical pseudopotential method, it can model the situation beyond thin cylinder limit [33]. It is also very interesting and valuable to search for possible non-abelian FCI phase in TBCB system. The problem about the coexistence of FCI phase and charge ordered phase [95] is also beyond one’s intuitive knowledge, which may occur in TBCB. Also, it is worth to find the connection between higher Chern number FCI and topological quantum chemistry in order to find material realization [96]. For spinful case there may be much richer competing phases, e.g., ferromagnetic phase. Also it is very interesting to study the spin-valley locking effect and the compitition between FCI and valley-polarized phase like $\rm{MoTe_{2}}$  [14] or spin-valley composite phase in moiré trilayer graphene [97, 98] case. The open questions mentioned above are worth further research in the future.
@@ Acknowledgements
We thank Jian Kang, Wei Li, Daniel E.Parker, Han-Qing Wu, Jia-Qi Cai, Gao-Pei Pan, Ming-Rui Li, Jun-Kai Dong, Ti-Xuan Tan, Zhongqing Guo, Wang-Qian Miao, Yunzhe Liu, Xiaohan Wan, Mu-Wei Wu for valuable discussions. J. Z. M. and D. X. Y. are supported by National Key Research and Development Program of China (2022YFA1402802), National Natural Science Foundation of China (92165204), Leading Talent Program of Guangdong Special Projects (201626003), Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (No. 2022B1212010008), and Shenzhen International Quantum Academy. J.-Y. C. is supported by National Natural Science Foundation of China (12304186), Open Research Fund Program of the State Key Laboratory of Low-Dimensional Quantum Physics (KF202207), Fundamental Research Funds for the Central Universities, Sun Yat-sen University (23qnpy60), Innovation Program for Quantum Science and Technology 2021ZD0302100, Guangzhou Basic and Applied Basic Research Foundation (2024A04J4264), the Guangdong Basic and Applied Basic Research Foundation (2024A1515013065). R.Z.H is supported by a postdoctoral fellowship from the Special Research Fund (BOF) of Ghent University.

### v2

@@ VI Conclusion and discussion
In this manuscript, we study a putative $C=2$ abelian FCI phase in $\nu=1/5,\kappa=0$ MATBCB. In the single particle level, we obtain the lowest topological flat bands with $C=\pm 2$. The band width is about $1.5\times 10^{-3}{\rm meV}$ at the first magic angle $\phi=1.608^{\circ}$. The almost uniform Berry curvature and tiny trace condition violation indicate the ideal quantum geometry condition to host FCI phase. In the many-body level, we first present the many body energy spectrum and spectrum flow for projected Coulomb interaction model of MATBCB. We find there is topological ground state manifold candidate with GSD=10. Next we examine the PES quasi-hole counting of the ground state manifold candidate. We find the counting meets the $(1,5)_{C=2}$ GPP as Halperin spin singlet. This counting is a smoking gun evidence of $C=2$ FCI. When the parameters are not ideal, such as the chiral ratio deviates from zero, there will be a phase transition which breaks the FCI phase. We estimate that $\kappa_{c}\approx 0.7050$. More comprehensive research on the FCI relevant phase transition will be left for future work.
However, one should still keep in mind that the results and arguments above are all built on the assumption that the translation symmetry is preserved. When the translation symmetry is broken, the internal SU(C) color symmetry can be entangled with spatial symmetry. Especially, the SU(C) symmetry and spatial translation symmetry will combine to form a generalized (pseudo) magnetic translation symmetry [33, 107]. In that case, we conjecture that the SU(C) color symmetry will also be broken accompanied with the translation symmetry broken. An example is the $\nu=1/6,C=2$ flat band system [70]. At that situation, Halperin color singlet quasi-hole counting ansatz based $(k,r)_{C}$ GPP may be violated. A more general pseudo potential method under $n_{x}-k_{y}$ basis has to be used [33, 8, 108]. Moreover, under $n_{x}-k_{y}$ basis it is convenient to simulate fractional filling system with numerical methods like DMRG [67, 66], which can handle larger size compared with ED. Compared with analytical pseudopotential method, it can model the situation beyond thin cylinder limit [33]. It is also very interesting and valuable to search for possible non-abelian FCI phase in TBCB system. The problem about the coexistence of FCI phase and charge ordered phase [92] is also beyond one’s intuitive knowledge, which may occur in TBCB. Also, it is worth to find the connection between higher Chern number FCI and topological quantum chemistry in order to find material realization [109]. For spinful case there may be much richer competing phases, e.g., ferromagnetic phase. Also it is very interesting to study the spin-valley locking effect and the compitition between FCI and valley-polarized phase like $\rm{MoTe_{2}}$  [14] or spin-valley composite phase in moiré trilayer graphene [110, 111] case. The open questions mentioned above are worth further research in the future.
@@ Acknowledgements
We thank Jian Kang, Wei Li, Daniel E.Parker, Han-Qing Wu, Jie Wang, Ying-Hai Wu, Jia-Qi Cai, Gao-Pei Pan, Ming-Rui Li, Jun-Kai Dong, Ti-Xuan Tan, Zhongqing Guo, Wang-Qian Miao, Yunzhe Liu, Xiaohan Wan, Mu-Wei Wu for valuable discussions. J. Z. M. and D. X. Y. are supported by National Key Research and Development Program of China (2022YFA1402802), National Natural Science Foundation of China (92165204), Leading Talent Program of Guangdong Special Projects (201626003), Guangdong Provincial Key Laboratory of Magnetoelectric Physics and Devices (Grant No. 656 2022B1212010008), Guangdong Fundamental Research Center for Magnetoelectric Physics, and Shenzhen International Quantum Academy. J.-Y. C. is supported by National Natural Science Foundation of China (12304186), Open Research Fund Program of the State Key Laboratory of Low-Dimensional Quantum Physics (KF202207), Fundamental Research Funds for the Central Universities, Sun Yat-sen University (23qnpy60), Innovation Program for Quantum Science and Technology 2021ZD0302100, Guangzhou Basic and Applied Basic Research Foundation (2024A04J4264), the Guangdong Basic and Applied Basic Research Foundation (2024A1515013065). R.Z.H is supported by a postdoctoral fellowship from the Special Research Fund (BOF) of Ghent University.